OpenStudy (anonymous):

When 75.5 grams of phosphorus pentachloride react with an excess of water, as shown in the unbalanced chemical equation below, how many moles of hydrochloric acid will be produced? Please show all your work for the calculations for full credit.

OpenStudy (chmvijay):

where is equation

OpenStudy (anonymous):

PCl5 + H2O yields H3PO4 + HCl

OpenStudy (chmvijay):

balance it

OpenStudy (anonymous):

this is what im asking u to do !!! this is the question

OpenStudy (chmvijay):

PCl5 +4H2O = H3PO4 + 5HCL

OpenStudy (chmvijay):

wt is molecular weight of PCl5

OpenStudy (anonymous):

we dont want any molecular mass , i just want to know "how many moles of hydrochloric acid will be produced? Please show all your work for the calculations for full credit. "

OpenStudy (chmvijay):

if u say molecular weight then i can claculate it

OpenStudy (anonymous):

75.5 grams

OpenStudy (chmvijay):

no molar mas of Pcl5 is not correct

OpenStudy (chmvijay):

208 gram of Pcl5 gives rise to = 5 moles of Hcl 75.5 grams of pcl5 gives rise to =(5*75.5)/208 = solve it this the answer

OpenStudy (anonymous):

is this the final answer 1.814

OpenStudy (chmvijay):

good in maths right answer 1.81